There are 385 surfers in Surf City. Two thirds
of the boys are surfers and one-twelfth of the girls
are surfers. If there are two girls for every boy, then how
many boys and how many girls are there in Surf City?

Answer by stanbon(75887)

There are 385 surfers in Surf City.
Two thirds of the boys are surfers and one-twelfth of the girls
are surfers.
If there are two girls for every boy, then how
many boys and how many girls are there in Surf City?
-----
Equations:
surfers + surfers = surfers
(2/3)b + (1/12)g = 385
g = 2b
-----------------------
Modify the 1st equation:
8b + g = 12*385
---
Substitute for "g" and solve for "b":
8b + 2b = 12*385
10b = 4620
b = 462 (# of boys)
-----
Since g = 2b, g = 924 (# of girls)
=========================================
Cheers,
Stan H.